OpenStudy (anonymous):

What is the simplified form of x minus 5 over x squared minus 3x minus 10 • x plus 4 over x squared plus x minus 12 ?

OpenStudy (anonymous):

do you even know how to do this?

OpenStudy (anonymous):

you have to factor them first

Nnesha (nnesha):

\[\huge\rm \frac{ x-5 }{ x^2 -3x -10 } \times \frac{ x+4 }{ x^2 +x -12 }\] like this right ?

OpenStudy (anonymous):

yes

Nnesha (nnesha):

yes right factor both denominator

Nnesha (nnesha):

do you know to factor quadratic equations ?

Nnesha (nnesha):

how*

OpenStudy (anonymous):

you find two mutilply numbers that equals -10 and add up to -3

Nnesha (nnesha):

yep^

OpenStudy (anonymous):

thats where i need help

Nnesha (nnesha):

start from 1 10 times 1 =10 5 times 2 = 10 these number you can try to factor now change signs and see which one give you -3 when you add or subtract

OpenStudy (anonymous):

-5 and 2

Nnesha (nnesha):

yep right so because leading coefficient is one so you can right like this (x -5) and x-2) now factor this x^2 +x -12

OpenStudy (anonymous):

-3 and 4

Nnesha (nnesha):

yep right \[\huge\rm \frac{ x-5 }{ (x-5)(x+2) } \times \frac{ x+4 }{ (x-3)(x+4) }\]

OpenStudy (anonymous):

now you just cross out the like terms

Nnesha (nnesha):

yep

OpenStudy (anonymous):

i get now but can you help me with one more please

Nnesha (nnesha):

sure :)

OpenStudy (anonymous):

After the fraction x plus 2 all over 3 minus the fraction x plus 1 all over 2x have been combined using the least common denominator of 6x, what is the numerator?

OpenStudy (anonymous):

what do u do???

Nnesha (nnesha):

is this ur question ?